Shift Summary for Owl, 6/22/99

Robert Tschirhart (tsch@fnal.gov)
Tue, 22 Jun 1999 07:03:05 -0500

Shift Summary for Owl shift, Jun. 22, 1999

On Shift: V. Jejer, B. Tschirhart
plus Elliott hanging on for a while.

Hours of beam delivered at all intensities: 7.00h
Hours of beam delivered at nominal intensity: ~5.00h
Hours of E832 data (special runs not included): 3.50h
Hours of special runs (laser scans, muon runs...): 0.00h

Accounting of lost time: ktev2 crash, DAQ problems, trigger timing probs.

========================================================================
Runs ended this shift:

Run 13682 - BEAM_MU_4E12
13683 - BEAM_MU_4E12.
13685-13688 Debugging Beam_832, started 13689 which is good run.

========================================================================

Accesses during shift:

========================================================================
Overview of shift:

Rough start. Inherited muon runs from evening shifts that had trigger
timing shifts of about 1 nsec as seen in hodoscope TDCs. At the start of
the owl shift ktev2 crashed, followed by lots of front-end trouble,
followed by a trigger timing adjustment, followed then finally by a
respite of smooth sailing.

---> +1nsec delay in Tevatron RF starting at run 13689. <---

========================================================================
Description of Shift:

Run Spill Time Action
------------------------------------------
00:30 ktev2 crashes, epicure terminal
freezes. ECC restarts and reboots
ktev2. No space in root directory
due to debris from previous crash
yesterday. ECC cleans up, now is ok.

00:40 P. Shanahan phones and says that
CsI channel 2310 should come out at
the next beam downtime. This channel
has blown comparator symptoms. No
beam downtime is in sight.

13684 We try to init a new run, but ktev2
crash wiped out EOF marks on the new
tape set that only has the two muon
runs. Since these muon runs are written
to disk, we proceed to OVERWRITE THESE
RUNS with the hoped for E832_BEAM run,
which is the increasingly elusive goal of
our shift.

02:30 MCR calls, and reports trouble maintaining
our intensity request. Typically we are
running at 2-4E12 as compared with the
nominal 5E12 request.

13685-> 01:15 -> Runs would not start due a fascinating
13688 02:50. variety of front-end problems with
Stream-5 and Stream-6 that were eventually
fixed with a reload of the HCC and Etot
trigger. Details are in the logbook.

13688 03:10 Finally got run 13688 up an running.
Delay of RF with respect to previous
running is the same. TDC distribution
of hodoscpes are all typically off by
1 nsec.

13689 03:20 Added 1 nsec to RF delay chain. TDC
distributions now look significantly
better. Additional delay added just
in front of the LRS-821 discriminator
in RR154.

03:30 Gas check.

05:01 Crate-53 hangups up again. Requires
a KFREND to fix.

06:40 Intensity fluctuations, varying from
2E12 to 4E12, BOA still looks good though.

07:00 Hand off to day shift.

========================================================================
Items to Follow up on Next Shift

* If downtime, notify Ed about opportunity to swap out CsI 2310.

*The SEM calibration will likely change today. Rick Ford/Rick Coleman
will notify us, expect about a 20% adjustment. We should strive to
stay at the same real intensity --> request will have to change after
adjustment.